OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation. y=4 when x=6 .

OpenStudy (saifoo.khan):

LEESHAAAH

OpenStudy (saifoo.khan):

y = k/x

OpenStudy (saifoo.khan):

4*6 = k k = 24

OpenStudy (saifoo.khan):

therefore: y = 24/x

OpenStudy (anonymous):

Thanks Saifoooo . (;

OpenStudy (saifoo.khan):

uw LEEE-SHAAAAHHH.

OpenStudy (anonymous):

- Do you know how to help w/ thee rest of my questions ?

OpenStudy (saifoo.khan):

i guess so..